OpenStudy (anonymous):

Write the equation of the line with the slope = -2/5 that passes through the point (-2, 6)

OpenStudy (anonymous):

Use point-slope form. y-y1=m(x-x1) Plug in what you have and solve it in terms of y.

OpenStudy (anonymous):

You have a point in the form (x1,y1) and a slope m so use point-slope form\[y - y1 = m(x-x1)\] then solve for y

OpenStudy (anonymous):

I did that but I come up with y=-2/5x+26/5. That is not one of the options for ABC or D

OpenStudy (anonymous):

whats your choices?

OpenStudy (anonymous):

Write the equation of the line with the slope = -2/5 that passes through the point (-2, 6) Select one: A. 2x – 5y = -26 B. 2x + 5y = 26 C. 5x + 2y = 2 D. 5x – 2y = -22

OpenStudy (anonymous):

Ohh i see, solve each of those for y and find the one that matches up with what we came up with above.

OpenStudy (anonymous):

I don't understand?

OpenStudy (anonymous):

they dont have their equations solved for y like you did \[y = -\frac{2}{5}x + \frac{26}{5}\] since the right side has a common denominator in both terms we can do this \[y = \frac{-2x + 26}{5}\] now multiply both sides by 5 to cancel out the 5's on the right hand side \[5y = -2x + 26\] move the term with the x over to the left hand side like they have it \[5y + 2x = 26\] is what we get

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You're welcome